OpenStudy (anonymous):

(Definite integral) x * e^(x^2) dx from 1 to 3

OpenStudy (anonymous):

u sub.

OpenStudy (anonymous):

Oh, yeah I had a question about it:

OpenStudy (helder_edwin):

u have \[\large \int xe^{x^2}\,dx \] try the substitution \(u=x^2\)

OpenStudy (anonymous):

go ahead

OpenStudy (anonymous):

Sorry for the confusion, I've worked it out already, but I was just wondering why it's 1/2 and not 1/4 (see my attached image):

OpenStudy (helder_edwin):

what u did is wrong!! @TJ91x

OpenStudy (anonymous):

Uh oh. :(

OpenStudy (anonymous):

you would have to leave the 1/2u inside the integrand and it would cancel with u = e^(x^2) \[\int\limits_1^3 xe^{x^2}\, dx = \frac{1}{2} \int\limits_{e}^{e^9} du\]

OpenStudy (anonymous):

ooops... i goofed.

OpenStudy (anonymous):

yeah, let u = x^2

OpenStudy (anonymous):

you don't have an extra e^(x^2) for du.

OpenStudy (anonymous):

Ohh, okay, I see. I'm trying the substitution with x^2 instead of e^x^2

OpenStudy (anonymous):

Oh boy, that worked out much better:

OpenStudy (anonymous):

there you go

OpenStudy (anonymous):

Thanks guys. :)

OpenStudy (anonymous):

you're welcome!

OpenStudy (helder_edwin):

if \(u=x^2\) then \(du=2x\,dx\) or \(du/2=x\,dx\) so \[\large \int_1^3xe^{x^2}\,dx=\int_1^3e^u\,\frac{du}{2}= \frac{1}{2}\int_1^3e^u\,du=\frac{1}{2}e^u\Biggr|_1^3= \frac{1}{2}e^{x^2}\Biggr|_1^3 \] \[\large =\frac{1}{2}(e^9-e) \]